OpenStudy (anonymous):

Given sec s = 5/3, where sin s < 0, and cot t = 1, where 0 < t < π/2, find sin(s − t) and cos(s + t). sin(s − t) = ? cos(s + t) = ?

OpenStudy (anonymous):

hellooooo anybody wanna help?

OpenStudy (irishboy123):

indeed where do you want to start? \(\cot t = 1\) so \(t = \frac{\pi}{4}\)??

OpenStudy (anonymous):

I mean I keep getting 2√ / 10 as the answer.

OpenStudy (anonymous):

oops I meant √2/ 10

OpenStudy (anonymous):

you need \(\sin(t)\), \(\cos(t)\) , \(\sin(s)\) and \(\cos(s)\)

OpenStudy (anonymous):

did you get those?

OpenStudy (anonymous):

\[\sec(s)=\frac{5}{3}\to \cos(s)=\frac{3}{5}\to \sin(s)=-\frac{4}{5}\]

OpenStudy (anonymous):

\[\cos(t)=1\to t=\frac{\pi}{4}\to \cos(t)=\sin(t)=\frac{\sqrt2}{2}\]

OpenStudy (anonymous):

\[\cos(s+t)=\cos(s)\cos(t)-\sin(s)\sin(t)\] you have all the numbers you need to compute that

OpenStudy (anonymous):

thank you! I'm working on it right now

OpenStudy (anonymous):

k